OpenStudy (anonymous):

4x^2+9y^2-8x-18y-23=0 find the foci

OpenStudy (xapproachesinfinity):

do a bit of a massage to the equation

OpenStudy (anonymous):

\[\frac{ (x-1)^2 }{ 9 }+\frac{ (y-1)^2 }{ 4 }\]

OpenStudy (xapproachesinfinity):

hmm so it is ellipse

OpenStudy (anonymous):

It is

OpenStudy (anonymous):

I hate these so much

OpenStudy (xapproachesinfinity):

(1,1) here is the center of this ellipse the distance from the two foci's are equal to 2a

OpenStudy (xapproachesinfinity):

too bad i gotta go for dinner

OpenStudy (anonymous):

now I am sht out of luck

Nnesha (nnesha):

it's a same question right fo find foci first you need c

Nnesha (nnesha):

to*

OpenStudy (anonymous):

It is pose to be (x1,y1),(x2,y2)

OpenStudy (anonymous):

written in that form :(

OpenStudy (anonymous):

open study lagging

Nnesha (nnesha):

\[\frac{ (x-1)^2 }{ 9 }+\frac{ (y-1)^2 }{ 4 }\] a^2 = 9 b^2 = 4 use this formula \[\huge\rm c^2 = a^2 - b^2\] solve for c and bec you already know it's horizontal that's why you should add c into x coordinate \[\huge\rm (h \pm c, k)\]

OpenStudy (anonymous):

c^2=65

Nnesha (nnesha):

\[\frac{ (x-1)^2 }{ 9 }+\frac{ (y-1)^2 }{ 4 }= 1\] don't forget one

Nnesha (nnesha):

hmm how did you get 65 ??

OpenStudy (anonymous):

nvm

Nnesha (nnesha):

equation  so a^2 = 9 and b^2 = 4